A confidence interval estimate is desired for the gain in a circuit on a semiconductor device. Assume the gain is normally distributed with standard deviation σ = 20.


a) Find a 95% confidence interval for μ when n = 10 and x̄ = 1000.


b) Find a 95% confidence interval for μ when n = 25 and x̄ = 1000.


c) Find a 99% confidence interval for μ when n = 10 and x̄ = 1000.


d) Find a 99% confidence interval for μ when n = 25 and x̄ = 1000.

1
Expert's answer
2022-04-27T00:42:36-0400

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The critical value for "\alpha = 0.01" is "z_c = z_{1-\alpha/2} =2.5758."

a)


"CI=(1000-1.96\times\dfrac{20}{\sqrt{10}}, 1000+1.96\times\dfrac{20}{\sqrt{10}})"

"=(987.604,1012.396)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "987.604 < \mu < 1012.396," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(987.604,1012.396)."


b)


"CI=(1000-1.96\times\dfrac{20}{\sqrt{25}}, 1000+1.96\times\dfrac{20}{\sqrt{25}})"

"=(992.16,1007.84)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "992.16 < \mu < 1007.84," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(992.16, 1007.84)."


c)


"CI=(1000-2.5758\times\dfrac{20}{\sqrt{10}}, 1000+2.5758\times\dfrac{20}{\sqrt{10}})"

"=(983.709,1016.291)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "983.709 < \mu < 1016.291," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(983.709,1016.291)."


d)


"CI=(1000-2.5758\times\dfrac{20}{\sqrt{25}}, 1000+2.5758\times\dfrac{20}{\sqrt{25}})"

"=(989.697,1010.303)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "989.697 < \mu < 1010.303," which indicates that we are 99% confident that the true population mean "\mu" is contained by the interval "(989.697,1010.303)."